OpenStudy (anonymous):

what equation represents the line that passes through the point (1,5) and has a slop of -2

OpenStudy (lieutenantgeneral):

first of all, use the gradient formula y2-y1/x2-x1 -1-5/7-1 -6/6= -1 use y-b=m(x-a) y-5=-1(x-1) y-5=-x+1 y=-x+6 y=6-x